A population consists of the numbers 5, 7, 9,




11, 12, and 15. List all possible sample size of




2 from this population and give the following:




•A. Determine the number of sets of all possible random samples that




can be drawn from the given population.




• B. List all possible samples and compute the mean of each sample.




• C. Construct the sampling distribution.




•D. Construct a histogram of the sampling distribution of the means.

1
Expert's answer
2022-03-25T14:04:38-0400

A) The number of possible samples is "C(6,2)=15"


B) All possible samples and the means


C) Sampling Distribution



"P(X=6)=P(X=7)=P(X=8.5)=P(X=9)=P(X=9.5)=P(X=11)=P(X=10.5)=P(X=12)=P(X=11.5)=P(X=13)=P(X=13.5)=\frac{1}{15}"


"P(X=8)=P(X=10)=\frac{2}{15}"

D) Histogram